# Path-based packing of icosahedral shells into multi-component aggregates

### v1

@@ Introduction
We have demonstrated how to geometrically construct multi-shell icosahedra and verified their energetic stability in various systems. Another important point is to understand how they grow dynamically in physical processes. To this end, we have performed molecular dynamics (MD) growth simulations [39], in which atoms are deposited on pre-formed clusters. This type of simulations has been used to interpret previous nanoparticle growth experiments [22, 39]. The results are shown in Fig. 3, D-H and in figs. S7-S11.
When depositing Rb atoms on a Na@${}_{13}@$Rb32 cluster (Fig. 3D), consisting of shells (0,0), (0,1) and (1,1), the growth begins with a spontaneous symmetry breaking to form the first chiral shell of class Ch1, i.e. shell (1,2) (or (2,1)), corresponding to the Na@${}_{13}@$Rb@${}_{32}@$Rb72 cluster. Rb atoms self-organise to grow a chiral shell on an achiral core. This symmetry breaking is specific to BG shells, and is well predicted by the path rules. In contrast, growth on MC shells continues without symmetry breaking if atoms of the same species are deposited [22]. As Rb atoms continue to be deposited, more Ch1 shells grow. To change class after the shell (1,2), it is necessary to deposit atoms larger than Rb. From Fig. 3B, it appears that Cs atoms have the right size and in fact, depositing Cs atoms on Na@${}_{13}@$Rb@${}_{32}@$Rb72 (Fig. 3E) results in a transition to the Ch2 class, as shells (2,2) and (2,3) form spontaneously.
@@
Further confirmation of path rule predictivity is provided in Fig. 3, F and H. In Fig. 3F, the size mismatch for the pairs AuNi, AuCo, AuFe, AgNi, AgCo, AgCu is compared with the optimal estimates obtained from Eq. 1 for the addition of one MC and one Ch1 shell on a MC core. The mismatch much better corresponds to Ch1 than to MC shells. For AuCo, AuFe, AuNi, AgNi and AgCo, Ch1 shells should grow on MC cores of 147 or 309 atoms ($k=3$ or 4), while for AgCu the core should be larger, of 561 atoms ($k=5$). This is confirmed by MD simulations (Fig. 3, F and G) in which growth proceeds spontaneously along the Ch1 class for Ag deposited on Ni147 or Cu561 Mackay cores. More results are in Section S3 of the supplementary material.
@@
In summary, the growth on top of icosahedral seeds naturally proceeds according to our rules for drawing paths in the hexagonal plane. At each stage of the growth, two possible steps (class-changing or class-conserving step) are possible; among them, the system spontaneously take the step that better fits the size mismatch between atoms of the pre-existing shell and those the growing one. If atoms of the same species are deposited, i.e. with zero mismatch, the step associated to the lower optimal mismatch is taken, which is always the class-conserving one. If one continues to deposit atoms of the same type, further and further shells belonging to the same chirality class are formed. On the other hand, the class-changing step always requires the deposition of atoms of a different species, with larger radius and with size mismatch close enough to the optimal one.
@@
We note that almost perfect shell-by-shell growth is achieved in all simulations due to the fast diffusion of deposited atoms on top of the close-packed shells [22].
Extension to anti-Mackay shells and outlook
@@
The mapping of icosahedral structures into paths can be extended to other cases. Here we establish the extension to the generalized anti-Mackay (AM) icosahedral shells of Fig. 4, A-C. AM shells are achiral and not close-packed, since they contain non-vertex particles with coordination lower than six. Each AM shell is identified by a couple of non-negative integers, which here we call $(p,q)$, determining the disposition of particles in the triangular facet of the shell (see Fig. 4A). Shells with $q=0$ have been described in the original work of Mackay [34], who proposed the possibility of adding to the fcc tetrahedra of the Mackay icosahedron one more shell in hexagonal close-packed (hcp) stacking. Multi-shell icosahedra terminated by AM shells of different types have been observed in confined aggregates of colloidal particles [23, 40, 25].
@@
AM shells can be packed by using the same rules described for CK ones. The same elementary steps in the hexagonal plane are allowed, but now one can decide whether to consider the CK or the AM version corresponding to the endpoint of the step. In this way a larger variety of icosahedra can be built. Again, the stability of these structures is provided by the size mismatch between particles in different shells. The optimal size mismatch for icosahedra with AM shells can be estimated by using the same geometric considerations made for CK shells (see Section S1.5 of the supplementary material). In Fig. 4D we compare the stability of all possible shells that can be put on top of a Mackay core, namely MC, Ch1 and AM1. The optimal mismatch of the AM1 shells is intermediate between those of the MC and Ch1 shells (see also fig. S3). The mismatch for adding a AM1 shell on a 147-atom Mackay core is close to the one of AgCu, so that AM shells should grow by depositing Ag atoms on Cu a Mackay core of this size. This is verified by the MD simulations of Fig. 4E, in which we observe the formation of an Ag $(1,3)^{\ast}$ shell. The growth proceeds quite differently from that of Fig. 3H: a second shell of AM2 type is formed, followed by a shell of BG type, i.e. the class is changed at each step. This behaviour is due to the fourfold adsorption sites on the surface of AM shells, which act as adatom traps and naturally leads to the formation of AM shells of higher class (see fig. S12 for details on the growth mechanism). We note that traps make the growing structure much less smooth than in the case of Fig. 3H.
@@
Our construction can be further generalized. Let us mention a few possibilities. First, the paths can begin at any point in the lattice, instead of (0,0), so that the shells enclose an empty volume. Multi-shell structures enclosing a cavity are relevant to biological systems [41, 42, 32], and have been observed in metal clusters [43]. Second, the non-equivalent sites of a shell can be decorated with different types of particles, while maintaining icosahedral symmetry. For example, the vertex atoms in the surface shell of a metal cluster can be of a different species than the other atoms, thus becoming isolated impurities embedded in a surface of a different material. This possibility is relevant to single-atom catalysis [44]. Moreover, the mapping into paths may be applied to other figures obtained by cutting and folding the hexagonal lattice, such as octahedra and tetrahedra, and to other lattices, including Archimedean lattices [17], that can better accommodate particles with non-spherically symmetric interactions. In summary, the mapping into paths is a powerful tool for the bottom-up design of chiral and achiral aggregates of atoms, colloids and complex molecules. The unusual geometries of these aggregates can be of interest in various fields, e.g. in catalysis, optics and synthetic biology.
@@ Acknowledgments
The authors acknowledge the project PRIN2022 PINENUT of the Italian Ministry of University and Research. The authors thank Giovanni Barcaro, El yakout El koraychy, Alessandro Fortunelli, Alberto Giacomello, Laurence D. Marks, Richard E. Palmer, Emanuele Panizon, Michele Parrinello, Cesare Roncaglia, Giulia Rossi, Manoj Settem, Erio Tosatti and Jeffrey R. Weeks for a critical reading of the manuscript.
Supplementary materials
Methods
Supplementary Text
Figs. S1 to S12
Tables S1 to S3
References (46-53)
[…]
Figure 2: Mapping icosahedra into paths. A Path of a Mackay icosahedron: $i=5$ shells are assembled on top of each other. B The three possible cases when choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$ at each step of the path. C Binding energy per atom for clusters made of a core with five MC shells plus a sixth shell, which is either MC or Ch1. Interactions are of Lennard-Jones (LJ) type (see Methods Section in the supplementary material). The particles of the sixth shell differ from those in the core only by their size. The energy is in units of the $\varepsilon$ of the LJ potential. The structures are shown in D. E Comparison of the optimal mismatch between steps $(0,k)\to(0,k+1)$ and $(0,k)\to(1,k)$, corresponding to additional MC and Ch1 shells on an MC core, respectively. The optimal mismatch of Eq. 1 (GEO values) is compared to Lennard-Jones (LJ) and Morse potential results. The Morse potential data are given for three values (4,5,6) of the parameter $\alpha$, that regulates the width of the potential well (see Methods Section in the supplementary material). F Comparison between steps $(1,k)\to(1,k+1)$ and $(1,k)\to(2,k)$, corresponding to additional Ch1 and Ch2 shells on a core containing $k$ MC and one Ch1 shells, respectively. The optimal mismatch of Eq. 1 (GEO) is compared to LJ data.
Figure 3: Chiral cluster structures and growth sequences. A A path connecting all BG shells through their nearby chiral shells. Black and grey arrows indicate paths with right and left-handed chiral shells. A generic path of this type may alternate both chiralities, corresponding to combinations of black and grey arrows. Shells are enumerated along the black-arrow path. B Optimal mismatch $sm_{i,i+1}$ according to Eq. 1 for the path in A. Points connected by brown and green lines correspond to changes of species every shell and every two shells. The optimal $sm_{i,i+1}$ are compared to the mismatch between pairs of atomic species, indicated by squares and diamonds. The triangles correspond to the optimal mismatch between two MC shells and two Ch1 shells. C Structures for $i=2,3,4$ along the path of A, with their compositions related to specific systems. D-E Growth sequences from MD simulations for alkali metal clusters and corresponding paths. D Rb atoms are deposited on an icosahedral Na@${}_{13}@$Rb32 seed, E the seed is Na@${}_{13}@$Rb104 on which Cs atoms are deposited. F Size mismatch for some transition metal pairs compared to the optimal mismatch for a MC and a Ch1 shell on a MC core (green and brown lines, respectively), as a function of core size. G-H Snapshots from growth simulations and corresponding paths for the deposition of Ag atoms on G Ni147 and H Cu561 cores. All simulation snapshots in D-E, G-H are taken at magic sizes for the corresponding paths.
[…]
@@ Methods
Methods
Construction of the structures. Icosahedral structures are built by assembling Caspar-Klug and anti-Mackay shells by a purposely written C++ code. The code takes as input the geometric features of each shell, i.e. the indexes $h$ and $k$ of the CK construction, the distance $d$ between particles, and the shell type (either CK or AM). For CK shells, the first 12 particles are placed in the vertices of the icosahedron of edge legth $\sin(2\pi/5)\sqrt{T}\,d$. The other particles are placed on the icosahedral facets according to the CK scheme: on each facet plane a 2D hexagonal lattice is built, which is rotated of an angle of amplitude $\theta_{CK}$ with respect to one of the facet edges; particles are placed on lattice points falling within the facet. For AM shells, the indexes $p$ and $q$ are calculated as $p=k-h+1$, $q=h-1$ if $k\geq h$, $p=h-k+1$, $q=k-1$ otherwise. The first 12 particles are placed in the vertices of the icosahedron of edge length $(p+\sqrt{3}\,q+\sqrt{3}-1)\,d$. The other particles are placed on the icosahedral facets according to scheme of Fig. 4A in the main text. A more detailed description of CK and AM shells and of the procedure for constructing them can be found in Section S1 of the Supplementary Text.
$E=\frac{1}{2}\sum_{i\neq j}u(r_{ij}),$ (3)
where the dimensionless parameter $\alpha$ regulates the width of the potential well, that decreases with increasing $\alpha$. For $\alpha=6$ both LJ and Morse potential have the same width of the well, i.e. the same curvature at the well bottom. In the simulations of Fig. 2, C, E and F in the main text all particles were given the same value of $\varepsilon$ (and of $\alpha$ for the Morse potential), but particles of the outer shell and of the core were given different sizes (i.e. $r_{m}=r_{m,c}$ for core particles and $r_{m}=r_{m,s}$ for outer shell particles). For interactions between particles of the core and the outer shell $r_{m,cs}=(r_{m,c}+r_{m,s})/2$. The structures were locally relaxed by quenched molecular dynamics [46] to reach the position of the local minimum in the energy landscape.
Molecular dynamics (MD) growth simulations. Molecular Dynamics (MD) growth simulations are made by molecular dynamics using the same type of procedure adopted in Ref. [39, 22]. The equations of motion are solved by the Velocity Verlet algorithm with a time step of 5 fs for the simulations of AgNi, AgCu, AgCo, AuCo, AuFe, AgPdNi and 2 fs for the simulations of NaK, NaRb, NaKRb, NaRbCs. In all simulations, the temperature is kept constant by an Andersen thermostat with a collision frequency of 5 1011 s-1. Simulations start from a seed, which is an initial cluster, then atoms are deposited one by one on top of it in an isotropic way from random directions at a constant rate. The simulation of Fig. 3D in the main text was started from a Na@${}_{13}@$K32 Bergman-type seed corresponding to the path arriving to $(h,k)=(1,1)$ and Rb atoms were deposited at a rate of 0.1 atoms/ns and at a temperature of 125 K. The simulation of Fig. 3E in the main text was started from a Na@${}_{13}@$Rb@${}_{32}@$Rb72 chiral seed corresponding to the path arriving to $(h,k)=(1,2)$ and Cs atoms were deposited at a rate of 0.1 atoms/ns at 125 K. The simulation of Fig. 3G in the main text was started from a Ni147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 0.1 atoms/ns at 450 K. The simulation of Fig. 3H in the main text was started from a Cu561 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,5)$ and Ag atoms were deposited at a rate of 0.1 atoms/ns at 450 K. The simulation of Fig. 4E in the main text was started from a Cu147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 1 atoms/ns at 350 K. For all systems, atom-atom interactions were modelled by an atomistic force field, which is known as Gupta potential [49]. Form and parameters of the potential can be found in Refs. [50, 51, 52, 53].
Supplementary Text
Appendix S1 Construction of the structures and optimal mismatch
@@
    the series in Fig. 3A in the main text, passing through shells of Bergman (BG) type each two steps; here we call it BG icosahedral series;
@@
    the series of core-shell icosahedra with a Mackay (MC) core surrounded by a thick layer of shells of chiral-1 (Ch1) type, as are obtained in the growth of Ni@Ag and Cu@Ag core@shell nanoparticles (see Figs. 3, G and H in the main text).
@@
The index $h$ is incremented by 1 every two steps, from 0 to $h_{max}$; the first and the second shell of index $h$ have $k=h$ and $k=h+1$, respectively. We consider an icosahedron made of $i$ shells. If $i$ is even, the maximum value of $h$ is
@@
Here we explain in details how the construction of CK shells is implemented in our C++ code. In the CK scheme, each icosahedral facet is an equilateral triangle whose vertices are lattice points of a 2D hexagonal lattice. The coordinates of the vertices are given with respect to the the primitive vectors $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$, forming an angle of $60^{\circ}$ (see Fig. SS1A): $V_{1}(0,0)$, $V_{2}(h,k)$, $V_{3}(-k,h+k)$, where the positive integers $h$ and $k$ are the indexes of the CK construction, that unambiguously determine the icosahedral shell structure. All lattice points within the triangle belong to the CK shell, i.e. a lattice point $P(i_{1},i_{2})$ belongs to the shell if $i_{1}$ and $i_{2}$ meet the following conditions:
@@
It is useful to write the point $P$ with respect to the vectors $\mathbf{a}_{1}^{\prime}$ and $\mathbf{a}_{2}^{\prime}$, lying on two adjacent sides of the triangle. These are obtained by rotating the primitive vectors of the hexagonal lattice counterclockwise of an angle $\theta_{CK}$ (see Fig. S1B). The primitive vectors can be written as
@@
Here we consider core-shell icosahedra of a different type. The internal core is a complete MC icosahedron, whereas the external shell is of Ch1 type. Again, we consider bi-elemental structures, in which the core and the shell are made of different particles, with ideal pair distances $d_{CORE}$ and $d_{SHELL}$. The CK indexes of the most external shell of the MC core are $(0,k)$, whereas the CK indexes of the shell are $(1,k)$: passing from the core to the shell we increment the index $h$ from 0 to 1, while the index $k$ is unchanged (here we follow the rules for building CK icosahedra that we have discussed in the main text). The radius of the core is

### v2

@@ 3 SHELLS 4 SHELLS
We have demonstrated how to geometrically construct multi-shell icosahedra and verified their energetic stability in various systems. Another important point is to understand how they grow dynamically in physical processes. To this end, we have performed molecular dynamics (MD) growth simulations [44], in which atoms are deposited on pre-formed clusters. This type of simulations has been used to interpret previous nanoparticle growth experiments [24, 44]. The results are shown in Figure 5 and in Supplementary Figures 7-11.
Figure 5: Growth sequences of chiral icosahedra. a Rb atoms are deposited on an icosahedral Na@${}_{13}@$Rb32 seed. b The seed is Na@${}_{13}@$Rb104, on which Cs atoms are deposited. c Size mismatch for some transition metal pairs compared to the optimal mismatch for a MC and a Ch1 shell on a MC core (green and brown lines, respectively), as a function of core size. d-e Snapshots from growth simulations and corresponding paths for the deposition of Ag atoms on d Ni147 and g Cu561 cores. All simulation snapshots in a-b, d-e are taken at magic sizes for the corresponding paths. In the top and bottom rows of the snapshot sequences we show the cluster surface and its cross section, respectively.
As a first case, we consider the growth of multi-element alkali clusters. In the previous Section we have demonstrated the stability of non-trivial multi-shell structures, which comprise BG and chiral shells in the same aggregate. Here we check whether these structure can be grown in a physically realistic process.
When starting from a Na@${}_{13}@$Rb32 cluster (Figure 5a), consisting of shells $(0,0)$, $(0,1)$ and $(1,1)$, our rules of Figure 2b predict that the next shell should be of Ch1 type, i.e. $(1,2)$ (or equivalently $(2,1)$), regardless of the size mismatch. Therefore, even when depositing atoms of the same type, we expect such symmetry breaking to take place. This is indeed the case, as shown in Figure 5a, in which in the first step of the growth a $(1,2)$ shell spontaneously form, leading to the Na@${}_{13}@$Rb@${}_{32}@$Rb72 cluster. The symmetry breaking upon deposition of the same species of atoms is specific to BG shells. In contrast, the growth on the other type of achiral shell, i.e. MC shell, continues without symmetry breaking if atoms of the same species are deposited [24].
At this point, according to our rules, two non-equivalent steps are possible: the class-conserving step to $(1,3)$ and the class-changing step to $(2,2)$. These steps correspond to different optimal mismatches, i.e. 0.06 and 0.11, respectively, as estimated by Eq. 1. Since we are keeping on depositing Rb atoms (i.e. with zero mismatch), we expect growth to proceed by the $(1,3)$ step, which has the lower optimal mismatch. Also in this case, our prediction is verified by the growth simulation of Figure 5a, which further continues within the Ch1 class.
To change class after the shell $(1,2)$, it is necessary to deposit atoms of bigger size than Rb. Specifically, the mismatch should be large enough to make the formation of the Ch1 $(1,3)$ shell unfavourable, thus addressing the growth towards the Ch2 class. At least, the mismatch between Rb and the deposited species should be larger than 0.06, which the optimal value for the formation of the Ch1 shell. From Figure 4b, it appears that Cs atoms have the right size (the mismatch between Rb and Cs is 0.08) and in fact, depositing Cs atoms on Na@${}_{13}@$Rb@${}_{32}@$Rb72 (Figure 5b) results in a transition to the Ch2 class, as shells $(2,2)$ and then $(2,3)$ form spontaneously during the growth.
@@
Further confirmation of path rule predictivity is provided in Figure 5c-e. In Figure 5c, the size mismatch for the pairs AuNi, AuCo, AuFe, AgNi, AgCo, AgCu is compared with the optimal estimates obtained from Eq. 1 for the addition of one MC and one Ch1 shell on a Mackay core. The mismatch much better corresponds to Ch1 than to MC shells. For AuCo, AuFe, AuNi, AgNi and AgCo, Ch1 shells should grow on Mackay cores of 147 or 309 atoms ($k=3$ or $4$), while for AgCu the core should be larger, of 561 atoms ($k=5$). This is confirmed by MD simulations (Figure 5d-e) in which a Ch1 shell spontaneously forms when depositing Ag atoms on Ni147 or Cu561 Mackay cores. The growth continues within the Ch1 class if further Ag atoms are deposited. More results are presented and discussed in Supplementary Note 3.
@@
In summary, the growth on top of icosahedral seeds naturally proceeds according to our rules for drawing paths in the hexagonal plane. At each stage of the growth, two possible steps (class-changing or class-conserving step) are possible; among them, the system spontaneously takes the step that better fits the size mismatch between atoms of the pre-existing shell and those of the growing one. If atoms of the same species are deposited, i.e. with zero mismatch, the step associated to the lower optimal mismatch is taken, which is always the class-conserving step. If one continues to deposit atoms of the same type, further and further shells belonging to the same chirality class are formed. On the other hand, the class-changing step always requires the deposition of atoms of a different species, with larger radius and with size mismatch close enough to the optimal one.
@@
Finally, we note that almost perfect shell-by-shell growth is achieved in all simulations due to the fast diffusion of deposited atoms on top of the close-packed shells [24].
Extension to anti-Mackay shells
@@
The mapping of icosahedral structures into paths can be extended to other cases. Here we establish the extension to the generalized anti-Mackay (AM) icosahedral shells of Figure 6. AM shells are achiral and not close-packed, since they contain non-vertex particles with coordination lower than six. Each AM shell is identified by a pair of non-negative integers, which here we call $(p,q)$, determining the disposition of particles in the triangular facet of the shell (see Figure 6a). Shells with $q=0$ have been described in the original work of Mackay [13], who proposed the possibility of adding to the fcc tetrahedra of the Mackay icosahedron one more shell in hexagonal close-packed (hcp) stacking. Multi-shell icosahedra terminated by AM shells of different types have been observed in confined aggregates of colloidal particles [25, 27, 45].
@@
[…]
Figure 7: Growth of generalized anti-Mackay structures. a Optimal mismatch for one AM1 shell on a Mackay core ($(0,k)\to(1,k)^{*}$) as a function of core size, compared with Ch1 ($(0,k)\to(1,k)$) and Mackay shells ($(0,k)\to(0,k+1)$), for Lennard-Jones clusters. GEO values are calculated by the formula derived in Supplementary Note 1.5. The mismatch between Ag and Cu is indicated. b,c Snapshots from MD growth simulations (with cluster surfaces and cross sections shown in top and bottom lines, respectively) and corresponding paths, with AM shells indicated by squares. In b Ag atoms are deposited on a Mackay Cu147 core. In c Pd atoms are deposited on a Ni core and then Ag atoms are deposited on the PdNi cluster. In the top and bottom rows of the snapshot sequences we show the cluster surface and its cross section, respectively.
@@
AM shells can be packed by using the same rules described for CK ones. The same elementary steps in the hexagonal plane are allowed, but now one can decide whether to consider the CK or the AM shell corresponding to the endpoint of the step. In this way a larger variety of icosahedra can be built. Again, the stability of these structures is ruled by the size mismatch between particles in different shells. The optimal size mismatch for icosahedra with AM shells can be estimated by using the same type of geometric considerations made for CK shells (see Supplemmentary Note 1.5).
In Figure 7a we compare the stability of all possible shells that can be put on top of a Mackay core, namely MC, Ch1 and AM1. The optimal mismatch of the AM1 shells is intermediate between those of the MC and Ch1 shells (see also Supplementary Figure 3). The mismatch for adding a AM1 shell on a 147-atom Mackay core is close to the one of AgCu, so that an AM shell should grow by depositing Ag atoms on Cu a Mackay core of this size. This is verified by the MD simulations of Figure 7b, in which we observe the formation of an Ag $(1,3)^{\ast}$ shell.
From there on, the growth proceeds in a different way from that of Figure 5e: a second shell of AM2 type is formed, followed by a shell of BG type, i.e. the class is changed at each step, even though atoms of the same type are deposited. This behaviour is due to the kinetic of the growth process. Specifically, it is due to fourfold adsorption sites on the surface of AM shells, which act as adatom traps and naturally lead to the formation of AM shells of higher class (see Supplementary Figure 11 for details on the growth mechanism). The fourfold traps make the growing structure much less smooth than in the case of Figure 5e. We note that, even though kinetic effects dominate the process, the growth proceeds step by step by incrementing only one shell index, as predicted by our rules.
Finally, in Figure 7c we show the growth of a ternary NiPdAg cluster. We start by depositing Pd atoms on a Ni147 Mackay cluster, that spontaneously arrange into an anti-Mackay $(1,3)^{\ast}$ shell. The mismatch between Ni and Pd is of 0.10, i.e. it is much larger than the optimal mismatch for growing a MC shell (0.04), and quite close to the optimal mismatch for a AM one (0.12, as estimated by geometrical considerations). On the other hand, the optimal mismatch for the formation of a Ch1 shell is quite larger (0.14, see Figure 7a). If the growth is continued by depositing Ag atoms, shells $(2,3)^{\ast}$ and then $(3,3)$ form. Such AM growth pathway appears to be unaffected by the mismatch, since it is observed also in the case of zero mismatch of Figure 7b. However, as for the CK shells, one can estimate the optimal size mismatch for changing the AM class, finding values sensibly larger than zero, and in the same range of those calculated for the CK shells. Therefore, depositing atoms of larger size, such as Ag (the mismatch between Pd and Ag is of 0.05), is expected to be beneficial for the stability of the resulting aggregate. Further deposition of Ag atoms leads to the formation of a chiral Ch3 shell, of type $(3,4)$. Indeed, the $(3,3)$ BG shell can be seen both as a generalised AM and a CK shell, and therefore it allows to obtain a very peculiar growth sequence, in which the two families of icosahedral shells are present in the same structure. We remark that the growth on top of a BG shell naturally leads to the formation of a chiral shell instead of an AM one, due to the lack of fourfold adsorption sites that are needed to form AM shells of class larger than 1.
Outlook
@@
Our construction can be further generalized. Let us mention a few possibilities. First, the paths can begin at any point in the lattice, instead of (0,0), so that the shells enclose an empty volume. Multi-shell structures enclosing a cavity are relevant to biological systems [34, 46, 47], and have been observed in metal clusters [41]. Second, the non-equivalent sites of a shell can be decorated with different types of particles, while maintaining icosahedral symmetry. For example, the vertex atoms in the surface shell of a metal cluster can be of a different species than the other atoms, thus becoming isolated impurities embedded in a surface of a different material. This possibility is relevant to single-atom catalysis [48]. Moreover, the mapping into paths may be applied to other figures obtained by cutting and folding the hexagonal lattice, such as octahedra and tetrahedra, and to other lattices, including Archimedean lattices [16], that can better accommodate particles with non-spherically symmetric interactions.
In summary, the mapping into paths is a powerful tool for the bottom-up design of chiral and achiral aggregates of atoms, colloids and complex molecules. The unusual geometries of these aggregates can be of interest in various fields, e.g. in catalysis, optics and synthetic biology.
Methods
Construction of the structures. Icosahedral structures are built by assembling Caspar-Klug and anti-Mackay shells by a purposely written C++ code. The code takes as input the geometric features of each shell, i.e. the indexes $h$ and $k$ of the CK construction, the distance $d$ between particles, and the shell type (either CK or AM). For CK shells, the first 12 particles are placed in the vertices of the icosahedron of edge legth $\sin(2\pi/5)\sqrt{T}\,d$. The other particles are placed on the icosahedral facets according to the CK scheme: on each facet plane a 2D hexagonal lattice is built, which is rotated of an angle of amplitude $\theta_{CK}$ with respect to one of the facet edges; particles are placed on lattice points falling within the facet. For AM shells, the indexes $p$ and $q$ are calculated as $p=k-h+1$, $q=h-1$ if $k\geq h$, $p=h-k+1$, $q=k-1$ otherwise. The first 12 particles are placed in the vertices of the icosahedron of edge length $(p+\sqrt{3}\,q+\sqrt{3}-1)\,d$. The other particles are placed on the icosahedral facets according to scheme of Figure 6a. A more detailed description of CK and AM shells and of the procedure for constructing them can be found in Supplementary Note 1.
$E=\frac{1}{2}\sum_{i\neq j}u(r_{ij}),$ (5)
where the dimensionless parameter $\alpha$ regulates the width of the potential well, that decreases with increasing $\alpha$. For $\alpha=6$ both LJ and Morse potential have the same width of the well, i.e. the same curvature at the well bottom. In the simulations of Figures 3a,c-d and 7a all particles were given the same value of $\varepsilon$ (and of $\alpha$ for the Morse potential), but particles of the outer shell and of the core were given different sizes (i.e. $r_{m}=r_{m,c}$ for core particles and $r_{m}=r_{m,s}$ for outer shell particles). For interactions between particles of the core and the outer shell $r_{m,cs}=(r_{m,c}+r_{m,s})/2$. The structures were locally relaxed by quenched molecular dynamics [49] to reach the position of the local minimum in the energy landscape.
Molecular dynamics (MD) growth simulations. Molecular Dynamics (MD) growth simulations are made by molecular dynamics using the same type of procedure adopted in Ref. [24, 44]. The equations of motion are solved by the Velocity Verlet algorithm with a time step of 5 fs for the simulations of AgNi, AgCu, AgCo, AuCo, AuFe, AgPdNi and 2 fs for the simulations of NaK, NaRb, NaKRb, NaRbCs. In all simulations, the temperature is kept constant by an Andersen thermostat with a collision frequency of 5$\times$1011 s-1. Simulations start from a seed, which is an initial cluster, then atoms are deposited one by one on top of it in an isotropic way from random directions at a constant rate. The simulation of Figure 5a in the main text was started from a Na@${}_{13}@$K32 Bergman-type seed corresponding to the path arriving to $(h,k)=(1,1)$ and Rb atoms were deposited at a rate of 0.1 atoms/ns and at a temperature of 125 K. The simulation of Figure 5b text was started from a Na@${}_{13}@$Rb@${}_{32}@$Rb72 chiral seed corresponding to the path arriving to $(h,k)=(1,2)$ and Cs atoms were deposited at a rate of 0.1 atoms/ns at 125 K. The simulation of Figure 5c was started from a Ni147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 0.1 atoms/ns at 450 K. The simulation of Figure 5d was started from a Cu561 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,5)$ and Ag atoms were deposited at a rate of 0.1 atoms/ns at 450 K. The simulation of Figure 7b was started from a Cu147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Ag atoms were deposited at a rate of 1 atoms/ns at 350 K. The simulation of Figure 7c were started from a Ni147 Mackay icosahedral seed corresponding to the path arriving to $(h,k)=(0,3)$ and Pd atoms were deposited at a rate of 0.1 atoms/ns at 400 K. Then Ag atoms were deposited on a Ni147@Pd132 seed terminated by a $(1,3)^{\ast}$ AM1 shell, at a rate of 0.1 atoms/ns at 300 K. For all systems, atom-atom interactions were modelled by an atomistic force field, which is known as Gupta potential [52]. Form and parameters of the potential can be found in Refs. [53, 54, 55, 56].
Data availability
The coordinates of the clusters shown in the figures are available upon request to the authors.
Code availability
The codes for the Molecular Dynamics simulations and for constructing the multi-shell icosahedra are available upon request to the authors.
@@ Acknowledgments
The authors acknowledge the project PRIN2022 PINENUT of the Italian Ministry of University and Research. The authors thank Giovanni Barcaro, El yakout El koraychy, Alessandro Fortunelli, Alberto Giacomello, Laurence D. Marks, Mauro Moglianetti, Richard E. Palmer, Emanuele Panizon, Michele Parrinello, Cesare Roncaglia, Giulia Rossi, Manoj Settem, Erio Tosatti and Jeffrey R. Weeks for a critical reading of the manuscript.
@@
Appendix S1 Construction of structures and optimal mismatch
@@
    the series in Figure 4a in the main text, passing through shells of Bergman (BG) type each two steps; here we call it BG icosahedral series;
@@
    the series of core-shell icosahedra with a Mackay (MC) core surrounded by a thick layer of shells of chiral-1 (Ch1) type, as are obtained in the growth of Ni@Ag and Cu@Ag core@shell nanoparticles (see Figure 5d-e in the main text).
@@
The index $h$ is incremented by $1$ every two steps, from 00 to $h_{max}$; the first and the second shell of index $h$ have $k=h$ and $k=h+1$, respectively. We consider an icosahedron made of $i$ shells. If $i$ is even, the maximum value of $h$ is
@@
Figure 8: Implementation of the CK construction.
Here we explain in details how the construction of CK shells is implemented in our C++ code. In the CK scheme, each icosahedral facet is an equilateral triangle whose vertices are lattice points of a 2D hexagonal lattice. The coordinates of the vertices are given with respect to the the primitive vectors $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$, forming an angle of $60^{\circ}$ (see Supplementary Figure 8a): $V_{1}=(0,0)$, $V_{2}=(h,k)$, $V_{3}=(-k,h+k)$, where the positive integers $h$ and $k$ are the indexes of the CK construction, that unambiguously determine the icosahedral shell structure. All lattice points within the triangle belong to the CK shell, i.e. a lattice point $P=(i_{1},i_{2})$ belongs to the shell if $i_{1}$ and $i_{2}$ meet the following conditions:
@@
It is useful to write the point $P$ with respect to the vectors $\mathbf{a}_{1}^{\prime}$ and $\mathbf{a}_{2}^{\prime}$, lying on two adjacent sides of the triangle. These are obtained by rotating the primitive vectors of the hexagonal lattice counterclockwise of an angle $\theta_{CK}$ (see Supplementary Figure 1b). The primitive vectors can be written as
@@
Here we consider core-shell icosahedra of a different type. The internal core is a complete MC icosahedron, whereas the external shell is of Ch1 type. Again, we consider bi-elemental structures, in which the core and the shell are made of different particles, with ideal pair distances $d_{CORE}$ and $d_{SHELL}$. The CK indexes of the most external shell of the MC core are $(0,k)$, whereas the CK indexes of the shell are $(1,k)$: passing from the core to the shell we increment the index $h$ from 00 to $1$, while the index $k$ is unchanged (here we follow the rules for building CK icosahedra that we have discussed in the main text). The radius of the core is